A vertical brass rod of circular section is loaded by placing a 5kg weight on top of it. If it's length is 50cm, it's radius of cross section 1cm and the young modulus of the material is 3.5power 19


1
Expert's answer
2021-10-27T08:48:01-0400

a) contraction in rod is given by

"\Delta l=\frac{mgl}{\pi r^2Y}"

Putting value of all variable and constants according to data of question

"\Delta l= \frac{5\times9.8\times50\times10^{-2}}{3.14\times1\times10^{-4}\times3.5\times10^{10}}"

"\Delta l=22.29\times10^{-7}m"


b) stored energy is given

"U=\frac{mg\Delta l}{2}"

"U=\frac{5\times9.8\times22.29\times10^{-7} }{2}"

"U=5.46\times10^{-5} J"